OpenStudy (anonymous):

Complete the square: (5/7)x^2+(9/10)x+(-13/70)=0

OpenStudy (anonymous):

ax^2+bx+c

OpenStudy (anonymous):

do i multiply by 7/5 to get the coefficient of the squared term to be 1?

OpenStudy (anonymous):

ax^2+2ab+b^2

OpenStudy (anonymous):

Divide everything by 5/7

OpenStudy (anonymous):

a=5/7 b=9/10

OpenStudy (anonymous):

firstly divide everything by 5/7 x^2 + (63/50)x-(-13/50) = 0 (x+63/100)^2 - 3969/10000 - 13/50 = 0 (x+63/100)^2 = 6569/10000 x = - (63/100)+-sqrt(6569/10000)

OpenStudy (anonymous):

then evaluate that .. that should hopefully be right (although i did it in my head, method is sound tho)

OpenStudy (anonymous):

looking over my numbers...

OpenStudy (anonymous):

got it! just went wrong on one of those huge fractions

OpenStudy (anonymous):

you are amazing!

OpenStudy (anonymous):

alternatively it is usually better instead of dividing it properly by 5/7, to put it outside a large bracket and leave it there, from start; 5/7[x^2+(63/65)x-(13/50) = 0

OpenStudy (anonymous):

forgot end bracket .. but u get the gist

OpenStudy (anonymous):

yes, great

OpenStudy (anonymous):

ok glad i could help

OpenStudy (anonymous):

I really appreciate it!